# Natural, Artificial, and Human Intelligences

### v1

@@ 4 The Word and Western Culture
A consideration of animal intelligence reveals differences with humans, but more so in terms of outcomes (such as science) rather than some essential ingredient of human intelligence, which is lacking in animals – one obvious exception is complex, written language. Debate on our origins is common throughout cultures. In some mythical traditions, humans are a kindred species with much to learn from wise animals; in others, humans are set apart, placed in the world by gods to preserve cosmic harmony or to provide service. The psychometric science has grown within a cultural tradition in which humans have always been set-apart by the faculties of word and reason: the defining characteristic of humanity has been assumed to be “the word made flesh”, a picture of humans which has persisted. Reviewing some of the roots of western science can help understand why chatbots present such an uncomfortable challenge, even if written language is not the whole story.
[…]
@@
Writing was elevated as a sacred medium in many civilizations. The Law of Moses (in the Jewish Torah, the first five books or Pentateuch of Christian Bibles) was meant to record the word of God. The Qur’an emphasizes the sacred status of its own words and respect for Jewish and Christian scriptures, whose followers are described as “people of the Book” (Surah 5:65–69). In China, the tradition of Confucian literary scholarship became central to the famous civil service examinations, in theory a great source of social mobility; these exams were entirely written (Elman,, 2013). Overall, there is evidence for an appreciation of the primacy of written language throughout human civilization, but there are exceptions. For example, in Sanskrit scholarship, Pāṇini’s Aṣṭādhyāyī (c. 5th to 4th century BCE), one of the earliest revered works on grammar and language, emphasized the sacred power of the Vedas (ancient Hindu scriptures), particularly through sound, rather than as static written texts.
During the scientific revolution, authors from Copernicus to Newton made efforts to separate their scientific, philosophical, and theological writings, but maintained the emphasis on human intelligence being uniquely rational and second only to that of God. Physical capabilities were considered less important, sometimes dismissed, most famously with Descartes’ “Je pense, donc je suis” (from Discours de la Méthode in 1637, the more famous Latin “Cogito ergo sum” appearing in Meditationes de Prima Philosophia in 1641). The term ‘scientific works’ became identified with what scientists had published, not physical objects. For example, in ancient times, Archimedes’ screw or Heron’s steam engine would be counted as scientific works. However, by the 1600s, Hooke’s work on mechanics and Leibniz’ on calculus would be taught as part of their scientific contributions, but not Hooke’s gear-cutting machines and Leibniz’ mine-pumps. This suggests an emerging distinction between science (expressed in writing) and technology, with the former considered a more intellectual activity.
Written science was enabled by key technology advances and ultimately led to more advanced scholarship. The printing press encouraged the spread of books and a preference for alphabetic and arithmetic symbols, instead of diagrams: for example, the Treviso Arithmetic, the first movable-type printed textbook of arithmetic, contains none of the diagrams used to visualize and compare ratios in a traditional copy of Euclid’s Elements of Geometry. Printing made it much easier to share accurate ‘ephemerides’ (tables of astronomical sightings), which, along with the mariner’s compass and the sextant, allowed more accurate long-distance navigation. The refracting telescope was carefully adapted for celestial accuracy by Galileo and Kepler, leading to the discoveries that planets had their own moons and elliptical orbits (American Institute of Physics,, 2025). Later, the desire to explain these patterns led to Newton’s axiomatic development of mechanics from three axioms, or laws, of motion.
Technological inventions gradually became more directly driven by scientific predictions. Literally, the pendulum swung — Galileo’s observation in the 1580s that the swing of a lantern kept the same period for large and small angles led to the pendulum clock built by Huygens in 1656 and a new era of temporal measurement accuracy, at least on land. Onboard ship, however, the pendulum’s period varied, and the naval chronometer, a clock accurate enough to keep time over several weeks, even as the time of noon onboard ship changed, was developed in the 1700s by John Harrison, in this case using careful craftsmanship, rather than scientific advances at the time.
@@
Maxwell’s electromagnetic theory in the 1860s went leaps further, explaining the propagation of light waves in traditional optics and predicting the behaviour of an entire spectrum of oscillations, invisible to human senses. During the 1880s, Hertz used these theoretical predictions to transmit and receive real radio waves. This formed a new relationship: science had brought about a technology that was beyond our natural imagination (Oerter,, 2006, Ch 1). In this new relationship, written works of science led directly to new kinds of machines, and scientific progress enabled a landslide of innovations throughout the late 1800s and early 1900s. Since that time, it has become assumed that written science is the main source of technological inspiration and reliability. Nonetheless, a small sample of inventions (fermentation, wheels, metalworking, textiles, clocks, gunpowder, printing, telescopes, steam engines, dried noodles, vacuum cleaners, velcro) demonstrates immediately that written science is just one path to invention (Mokyr,, 1992).
@@
Beyond science, the identification of human intelligence with written language skills has sometimes been even more emphatic. While the French Baccalaureate and German Abitur exams, formalized since the early-1800s, included practical and oral components, British qualifications, like the Chinese civil exams long before, were standardized entirely by writing. Given the emphasis on STEM education today, it is easy to assume that science education was developed along with science itself, but this would be an anachronism: Newton, Priestley, Cavendish, and Darwin did not go to secondary schools with classes in physics, chemistry, and biology, but instead, aside from geometry and algebra, a principal focus would have been classical languages and literature. British “grammar schools” had existed since the late middle-ages and the redirection of endowment funds for “other useful Branches of Literature and Science in addition to or (subject to the Provisions hereinafter contained) in lieu of the Greek and Latin Languages” was controversial enough to motivate the Grammar Schools Act of 1840 (Parliament of the United Kingdom,, 1840). The use of written examinations as the measure of intelligence was explicitly promoted in the Northcote and Trevelyan report that led to the British Civil Service exams. Their main stated goals included promoting merit over patronage, and also “establishing a proper distinction between mechanical and intellectual labour” (Northcote et al.,, 1854). While moral and physical fitness was to be vouched for in-person, for the purposes of choosing civil servants, intellectual labour was identified with writing. Overall, such tests prioritized written over manual skills: to fix a broken boiler, the person who could write down the problem, order the replacement parts etc., was considered intelligent, whereas the person who actually mended the boiler less so.
@@
Well before the time AI was invented, assumptions about the relationship between science, schooling, written words, and intelligence were part of the fabric of western European culture. For centuries, there were serious philosophical and scientific claims that abstract reasoning is impossible without language, and only recently, brain imaging studies have shown that these activities are dissociated from one another (Fedorenko et al.,, 2024). Language abilities have clearly helped to distinguish us from other species, but idolizing these abilities made us particularly vulnerable to the shock of fluent chatbots.
@@ 5 Artificial Intelligence Joins the Stage
The theory of computing was developed by Church and Turing in the 1930s, and Turing’s calculating machines were built in the hidden limelight of wartime codebreaking in the 1940s. By the end of the decade, Turing, von Neumann, and many collaborators had developed reprogrammable machines, whose instructions and calculations were stored in electronic memory and were capable of performing thousands arithmetic calculations within seconds, using cascading patterns of Boolean logic to process numbers represented in binary digits.
Shannon’s landmark Mathematical Theory of Communication (Shannon,, 1948) developed the quantification of information, using the statistics of character and word sequences in English as grounding examples. These are still foundational examples in statistical natural language processing: Shannon’s zeroth, first, and second-order approximations to English would nowadays be called $n$-gram models (Manning and Schütze,, 1999, Ch 6). Rather than trying to program and run these models on an early electronic computer, Shannon’s samples were taken from random pages in books – but he still accurately described the computational possibilities, and scaling limitations, of $n$-gram models.
@@
Turing’s paper on Computing Machinery and Intelligence in 1950 (Turing,, 1950) is a foundation of the field that was eventually called AI. Turing’s proposal was particularly important for text processing, introducing a challenge called the Imitation Game. Two humans, or a human and a machine, try to imitate one another, communicating directly with a judge, with interactions restricted to the inputs and outputs of a teleprinter. Whether the judge can distinguish a machine from a human through this interface alone has come to be called the Turing Test.
@@
Using the same distinction between physical and intellectual capacities as (Northcote et al.,, 1854), Turing reckoned that “The new problem has the advantage of drawing a fairly sharp line between the physical and the intellectual capacities of a man.” Turing carefully avoided the question of whether such a machine is ‘thinking’, saying instead that such behaviour would typically be regarded as demonstrating intelligence. His paper examined common objections to the idea that such a machine could be called intelligent (e.g., not having a soul or just learning ‘parrot-fashion’), arguing that they are rooted in prejudice and other mistakes, and speculated that the most effective way to build such a machine would be to create a ‘child machine’ with the programming and resources to be able to learn, and provide this machine with training experiences. He estimated that such a machine could be built by the year 2000 and, while that timeline was not achieved, Turing’s paper broadly anticipates the current approach to language processing, and the tendencies to resist these achievements, noting that this urge is “likely to be quite strong in intellectual people, since they value the power of thinking more highly than others, and are more inclined to base their belief in the superiority of Man on this power.”(Turing,, 1950, §6)
@@
Some of the core components of today’s chatbots were pioneered in the 1950s and 1960s. Most famously, Rosenblatt’s invention of the perceptron (Rosenblatt,, 1958) and Widrow and Hoff’s use of continuous, gradient-based iterations to train perceptron networks (Widrow and Hoff,, 1960) are generally considered the precursors of today’s neural networks (Raschka et al.,, 2022, Ch 2). Representing words as vectors learned from data, usually called ‘word embeddings’ since the success of the word3vec package in 2013 Mikolov et al., (2013), was pioneered in the SMART (System for the Mechanical Analysis and Retrieval of Text) project, led by Gerard Salton during the 1960s (Salton,, 1971). The success of the SMART system was demonstrated using early statistical evaluations of precision and recall, developed during the Cranfield experiments, whose purpose was indexing and retrieval techniques, particularly for scientific reference and military intelligence gathering (Cleverdon et al.,, 1966). The Cranfield project was foundational in information retrieval, not only for its results, but for establishing the practice of evaluating search engines by measuring the precision and recall of sets of search results, when compared with held-out test data. This practice was followed in the 1990s in the Text Retrieval Evaluation Conference series (Harman,, 1993), and submitting results to a “TREC Style Evaluation” is a key forerunner of many machine learning metrics and leaderboards, such as Kaggle and the MLCommons, used to compare models today. One of the key surprises from the Cranfield experiments was that automated indexing strategies, based purely on the words occurring in each document, outperformed indexing by human experts. This surprised, and initially dismayed researchers: “the results which seem to offend against every canon on which we were trained as librarians” (Cleverdon et al.,, 1966). ‘Computer’ had been a job title for skilled arithmeticians during World War II, but since the rules of arithmetic had been formalized, it stood to reason that a faster electronic machine that performed arithmetic could produce results that were just as good. ‘Indexer’, however, was a different matter: the job had been performed by expert librarians since the 1800s, and the notion that skilled work with words could be done by a machine just as well, if not better, was a shock. Statistical term-weighting (i.e., how much weight should a term be given in a retrieval engine) is an example: instead of weighting terms using controlled vocabularies of keywords and subject-headings, it worked better to assign weights, or importance scores, to each word in a document, based on how many times that word occurs in that document and how few other documents it occurs in. This family of weighting schemes is called tf-idf (term frequency, inverse document frequency). The insight that statistical term-weighting (or term specificity) is modelled so effectively by tf-idf is due particularly to (Sparck Jones,, 1972), and though it might be considered a blunt challenge to more formal logical semantics, it is today common to trace the linguistic and philosophical heritage of such statistical observations to the works of Firth, (1957) and Wittgenstein, (1953).
@@
Given such strong foundations, it may be surprising that statistical language processing using neural networks took so long to fully flourish. One reason is that it required decades of hardware advancements, including the development of graphics processing units and their adaptation to neural networks, to realize the potential of neural network methods. Another factor is that most research in AI and computational linguistics during the later decades of the 1900s was focused elsewhere, particularly on symbolic, logical, and syntactic methods. In linguistics, Chomsky wielded great influence and his early work on syntactic structures claimed both that “The most that can be reasonably be expected of linguistic theory is that it shall provide an evaluation procedure for grammars” and “probabilistic models give no particular insight into some of the basic problems of syntactic structure” (Chomsky,, 1957). In AI, the book Perceptrons: An Introduction to Computational Geometry (Minsky and Papert,, 1969), demonstrated key problems encountered by neural networks with the learning of non-additive functions, including cases as simple as the logical XOR. As key inventors of early neural networks, Minsky and Papert emphasized in their book that their objections were not meant to divide the field, but rather that research on neural networks would fade as other ideas took over; the 1988 edition included an epilogue saying “In any case, the 1970s became the golden age of a new field of research into the representation of knowledge.” The new techniques listed as part of this ’golden age’ included word-expert parsers, nonmonotonic logic, and planning procedures, which led many researchers to “set aside their interest in the study of learning in favor of examining the representation of knowledge.” With hindsight, this is seen more as a dead-end than a golden age and the beginning of a period that is sometimes called the “AI winter”. (Ironically, the term AI winter was also coined by Minsky, during the 1980s, as a warning against what might happen if expert systems did not bear fruit.) But for some decades, symbolic logic was considered to be the core of artificial intelligence and linguistic theory (Jones,, 1994) and the most established textbooks on mathematical and computational linguistics from this era focus on discrete symbolisms, barely mentioning statistics, probability, or continuous mathematical foundations such as vectors and gradients (Partee et al.,, 1993; Allen,, 1995).
@@
Today’s chatbots come from the very techniques that were supposed to have been left behind: machine learning in general and neural networks in particular are the cornerstones of AI today. During the 1980s, researchers including Hinton and LeCun developed the backpropagation method, which enables deep neural networks to be trained efficiently, by passing feedback on classification results back through layers in the network, using the chain rule from multivariable calculus (Rumelhart et al.,, 1986). Neural nets built in this way started to outperform expert systems at several tasks, for example image recognition for handwritten digits (LeCun et al.,, 1995). In 1997, Deep Blue, a chess computer program invented by Hsu, a doctoral student at Carnegie Mellon University, defeated the world chess champion Gary Kasparov, relying on a massively parallel expert fuzzy-rule-based system and effective use of a dataset of Grandmaster games (Campbell et al.,, 2002). In 2000s, the first neural net model for next-word prediction was trained successfully to produce relatively consistent natural language phrases, over longer context distances than those achieved with $n$-gram models. AI was beginning to produce systems that outperformed humans on complex symbolic tasks and, rather than doing this using rules created by human domain experts, this was achieved by leveraging fast massive computation. The game of Go, more combinatorically demanding than chess by a wide margin, was mastered by the Alpha Go system, using databases of real games for training, and then even more strikingly, by Alpha Go Zero, which learned solely by playing against itself, with no human supervision except for being given the rules and the resources to simulate play (Silver et al.,, 2017).
@@
Early on, attempts at AI development focused on creating systems which would excel at specific tasks, assuming that such systems put together would create a human-like, general intelligence: for example, Shannon famously proposed that the game of chess would be an ideal challenge for developing intelligent systems, since it is generally considered to require “thinking” for skilful play (Shannon,, 1950). Based on the combinatoric explosion of possible chess positions, Shannon argued that a brute-force approach, leveraging large computational power, would at best enable the creation of a weak chess player. But, fast forward to a few decades later and it turns out that one can design an extremely effective chess system, mainly by applying brute-force computation over large datasets. This is both an insight and a challenge to humans; some of the most poignant reflections on this theme have been written by Kasparov himself (Kasparov,, 2017).
@@
AI agents have become better by carefully leveraging larger example datasets and computational resources, a process which contrasts with the way humans accomplish similar expertise, typically requiring years of devotion Gladwell, (2008). Understandably, this is often unwelcome news to human experts, and the theme that general methods that leverage computation are ultimately the most effective, and by a large margin has even been dubbed “The Bitter Lesson” (Sutton,, 2019). But, the two questions, are we trying to construct systems which behave intelligently and are we trying to re-create human intelligence, are not the same, though the recognition of this difference has taken some time (Russell and Norvig,, 2022, Ch 1).
@@
The large language models (LLMs) that drive modern chatbots are a striking example of this engineering approach. They are built using neural networks to predict sequences of word-embedding vectors, the most recent key ingredient being the use of attention mechanisms, which enable sequential behaviour to be modelled using parallel computation (Vaswani et al.,, 2017). Subsequent models, including GPT-3 (Brown et al.,, 2020), Llama (Touvron et al.,, 2023), and DeepSeek (Liu et al.,, 2024), are trained and optimized to whole new levels of performance, though the core components of word embeddings, feed-forward neural networks, and attention-heads using projection (with the whole system trained to optimize a loss-function throughout the network using backpropagation) are fundamentally those established in the transformer model of Vaswani et al., (2017).
@@
These models use none of the syntactic parsers, logical forms, or knowledge representation techniques that dominated much of late-1900s AI. Instead, they work much as (Turing,, 1950) suggested: a ‘child machine’ is initialized with enormous computational resources and relatively few built-in procedures, just enough to learn from masses of training examples, after which it performs well — so well that its output, as exemplified by modern chatbots, produces language successfully enough that it would be considered a clear sign of intelligence in many established tests.

### v2

@@ 4 The Word and Western Culture
Electronic computing and artificial intelligence are recent products, driven by relationships between technology and science, that arose along with industrialization, in a culture that particularly prized written forms of intelligence. This section summarizes aspects of these cultural developments, and how they influenced some of the development and attitudes towards artificial intelligence.
In standard historic accounts, western civilization is the product of Hebrew religion and Greek science, becoming the Judeo-Christian tradition. Both of these roots embody the distinctiveness of human origins and intelligence, with an emphasis on knowledge expressed in words. In Genesis (Ch 3), Adam and Eve eat the fruit of “the tree of the knowledge of good and evil” and become self-aware. In Aristotle’s Politics (Bk VII Ch 3), “Virtue arises through nature (physis), habit (ethos), and reason (logos)”, and while some animals demonstrate ethos as well as physis, only humans have logos. This combination became central to Christianity, whose most mystical Gospel starts with “In the beginning was the Word” (Εν αρχην ο λογος, John Ch 1).
This did not originally imply that knowledge and intelligence are expressed in written words. Many great scholarly traditions have emphasized oral learning, and many scriptural and epic works were known and taught long before they were written down in canonical manuscripts. Plato’s Phaedrus dialogue raises fears that those who rely on writing “will not use their memories …they will appear to be omniscient and will generally know nothing” plato1953phaedrus, analogous to arguments today that artificial language technologies reduce human skills.
As Christianity became established, so did the belief that humans are verbal, rational animals, and that this is the divine core of our identity. Beyond religious education, this was taught in introductory logic and classification (we use classification in the logical sense) from the third century onwards, using the tree of Porphyry (Figure 1), perhaps history’s most famous decision-tree. It was used until the mid-1800s, when Boolean propositional calculus and truth-tables supplanted Aristotelean classification and syllogism in the teaching of logic (boole1854laws). The claim that humans are uniquely rational was not just part of biology and theology, it was taught as a foundational principle in western logic!
@@
The primacy of written words was established gradually, and not only in Europe. The Qur’an emphasizes the sacred status of its own words and respect for Jewish and Christian scriptures, whose followers are described as “people of the Book” (Surah 5:65–69). In China, the tradition of Confucian literary scholarship became central to the famous civil service examinations, and by the Ming period (1368–1644) had become the main route to official status. In theory, this was a great source of social mobility, and historically the first great example of ‘standardized testing’ (elman2013civil).
North-western Europe came to rely especially on writing as a source of truth and intelligence. The widespread use of printing, and the Protestant Reformation, emphasized the view that the Bible itself was literally the Word of God, rather than the more mystical ‘Word made Flesh’. In the Renaissance period, famous scholars were often multi-disciplinary and their works included sculpture, art, poetry and music; but for scholars of the later Enlightenment period, by ‘collected works’, we tend to mean a book with just writings. In music, a Western European tradition developed in which composers are celebrated, but by their ‘works’ we mean written manuscripts. (In Western classical ensembles, the paper with the written notation is even called the music, whereas in jazz ensembles, the paper documents are called charts.)
The independence of science itself from roots in natural philosophy and common engineering practices also evolved during this era. In the 1600s, the slide rule (based on Napier’s logarithms) and the pendulum clock (based on Galileo’s mechanics) were the first examples of theoretical discoveries, described in writing, leading directly to the creation of new devices.
@@
Maxwell’s electromagnetic theory in the 1860s went leaps further, explaining the propagation of light waves in traditional optics and predicting the behaviour of an entire spectrum of oscillations, invisible to human senses. During the 1880s, Hertz used these theoretical predictions to transmit and receive real radio waves. This formed a new relationship: science had brought about a technology that was beyond our natural imagination (oerter2006theory, Ch 1). In this new relationship, written works of science led directly to new kinds of machines, and scientific progress enabled a landslide of innovations throughout the late 1800s and early 1900s.
Since that time, it has often been assumed that written science is the main source of technological inspiration and reliability. Electronic computing is a clear example of technology that arose as the brainchild of written science, and has inherited some of these presuppositions (with many very successful results). Nonetheless, a small sample of inventions (fermentation, wheels, metalworking, textiles, clocks, gunpowder, printing, telescopes, steam engines, dried noodles, vacuum cleaners, velcro) demonstrate immediately that written science is just one path to invention (mokyr1992lever).
@@
Anglo-American educational practices have particularly emphasized written assessments. While the French Baccalaureate and German Abitur exams, formalized since the early-1800s, included practical and oral components, British qualifications, like the Chinese civil exams long before, were standardized entirely by writing. Given the emphasis on STEM education today, it is easy to assume that science education was developed along with science itself, but this would be an anachronism: Newton, Priestley, Cavendish, and Darwin did not go to secondary schools with classes in physics, chemistry, and biology, but instead, aside from geometry and algebra, a principal focus would have been classical languages and literature. British “grammar schools” had existed since the late middle-ages and the redirection of endowment funds for “other useful Branches of Literature and Science in addition to or (subject to the Provisions hereinafter contained) in lieu of the Greek and Latin Languages” was controversial enough to motivate the Grammar Schools Act of 1840 (grammarSchoolsAct1840). The use of written examinations as the measure of intelligence was explicitly promoted in the Northcote and Trevelyan report that led to the British Civil Service exams. Their main stated goals included promoting merit over patronage, and also “establishing a proper distinction between mechanical and intellectual labour” (northcote1854report). While moral and physical fitness was to be vouched for in-person, for the purposes of choosing civil servants, intellectual labour was identified with writing. Overall, such tests prioritized written over manual skills: to fix a broken boiler, the person who could write down the problem, order the replacement parts etc., was considered intelligent, whereas the person who actually mended the boiler less so.
@@
Well before the time the term AI was coined, assumptions about the relationship between science, schooling, written words, and intelligence were part of the fabric of western European culture. For centuries, there were serious philosophical and scientific claims that abstract reasoning is impossible without language, and only recently, brain imaging studies have shown that these activities are dissociated from one another (fedorenko2024language).
Language and reasoning abilities have clearly helped to distinguish us from other species, and writing as a technology has worked as a great catalyst for improving, and even helping invent, other technologies. Not surprisingly, the most successful industrial societies have prized and cultivated writing skills. However, such societies are prone to find the idea of a ‘thinking machine’ threatening. Turing even anticipated such cultural resistance, saying “It is likely to be quite strong in intellectual people, since they value the power of thinking more highly than others (turing1950computing, §6.II)”.
@@ 5 Artificial Intelligence Joins the Stage
The first electronic computers are a great example of a scientific discovery, based on written work, leading directly to new technologies. The theory of computing was developed by Church and Turing in the 1930s, and Turing’s calculating machines were built in the hidden limelight of wartime codebreaking in the 1940s. By the end of the decade, Turing, von Neumann, and many collaborators had developed reprogrammable machines, whose instructions and calculations were stored in electronic memory and were capable of performing thousands arithmetic calculations within seconds, using cascading patterns of Boolean logic to process numbers represented in binary digits.
@@
Turing’s paper on Computing Machinery and Intelligence in 1950 (turing1950computing) is a foundation of the field that was eventually called AI. Turing’s proposal was particularly important for text processing, introducing a challenge called the Imitation Game. Two humans, or a human and a machine, try to imitate one another, communicating directly with a judge, with interactions restricted to the inputs and outputs of a teleprinter. Whether the judge can distinguish a machine from a human through this interface alone has come to be called the Turing Test.
@@
Using the same distinction between physical and intellectual capacities as (northcote1854report), Turing reckoned that “The new problem has the advantage of drawing a fairly sharp line between the physical and the intellectual capacities of a man.” Turing carefully avoided the question of whether such a machine is ‘thinking’, saying instead that such behaviour would typically be regarded as demonstrating intelligence. His paper examined common objections to the idea that such a machine could be called intelligent (e.g., not having a soul or just learning ‘parrot-fashion’), arguing that they are rooted in prejudice and other mistakes, and speculated that the most effective way to build such a machine would be to create a ‘child machine’ with the programming and resources to be able to learn, and provide this machine with training experiences. He estimated that such a machine could be built by the year 2000 and, while that timeline was not achieved, Turing’s paper broadly anticipates the current approach to language processing, and some of the likely objections to ascribing intelligence to such machines.
@@
Some of the core components of today’s chatbots were pioneered in the 1950s and 1960s. Most famously, Rosenblatt’s invention of the perceptron (rosenblatt1958perceptron) and Widrow and Hoff’s use of continuous, gradient-based iterations to train perceptron networks (widrow1960adaptive) are generally considered the precursors of today’s neural networks (raschka2022machine, Ch 2). Representing words as vectors learned from data, usually called ‘word embeddings’ since the success of the word3vec package in 2013 mikolov2013efficient, was pioneered in the SMART (System for the Mechanical Analysis and Retrieval of Text) project, led by Gerard Salton during the 1960s (salton1971smart). The success of the SMART system was demonstrated using early statistical evaluations of precision and recall, developed during the Cranfield experiments, whose purpose was indexing and retrieval techniques, particularly for scientific reference and military intelligence gathering (cleverdon1966factors). The Cranfield project was foundational in information retrieval, not only for its results, but also for establishing the practice of evaluating search engines by measuring the precision and recall of sets of search results, when compared with held-out test data. This practice was followed in the 1990s in the Text Retrieval Evaluation Conference series (harman1993first), and submitting results to a “TREC Style Evaluation” is a key forerunner of many machine learning metrics and leaderboards, such as Kaggle and the MLCommons, used to compare models today.
One of the key surprises from the Cranfield experiments was that automated indexing strategies, based purely on the words occurring in each document, outperformed indexing by human experts. This surprised, and initially dismayed researchers: “the results which seem to offend against every canon on which we were trained as librarians” (cleverdon1966factors). ‘Computer’ had been a job title for skilled arithmeticians during World War II, but since the rules of arithmetic had been formalized, it stood to reason that a faster electronic machine that performed arithmetic could produce results that were just as good. ‘Indexer’, however, was a different matter: the job had been performed by expert librarians since the 1800s, and the notion that skilled work with words could be done by a machine just as well, if not better, was a shock.
@@
Given such strong foundations, it may be surprising that statistical language processing using neural networks took so long to fully flourish. One reason is that it required decades of hardware advancements, including the development of graphics processing units and their adaptation to neural networks, to realize the potential of neural network methods. Another factor is that most research in AI and computational linguistics during the later decades of the 1900s was focused elsewhere, particularly on symbolic, logical, and syntactic methods. In linguistics, Chomsky wielded great influence and his early work on syntactic structures claimed both that “The most that can be reasonably be expected of linguistic theory is that it shall provide an evaluation procedure for grammars” and “probabilistic models give no particular insight into some of the basic problems of syntactic structure” (chomsky1957syntactic). In AI, the book Perceptrons: An Introduction to Computational Geometry (minsky2017perceptrons), demonstrated key problems encountered by neural networks with the learning of non-additive functions, including cases as simple as the logical XOR. As key inventors of early neural networks, Minsky and Papert emphasized in their book that their objections were not meant to divide the field, but rather that research on neural networks would fade as other ideas took over; the 1988 edition included an epilogue saying “In any case, the 1970s became the golden age of a new field of research into the representation of knowledge.” The new techniques listed as part of this ‘golden age’ included word-expert parsers, nonmonotonic logic, and planning procedures, which led many researchers to “set aside their interest in the study of learning in favor of examining the representation of knowledge.” With hindsight, this is seen more as a dead-end than a golden age and the beginning of a period that is sometimes called the “AI winter”. (Ironically, the term AI winter was also coined by Minsky, during the 1980s, as a warning against what might happen if expert systems did not bear fruit.) But for some decades, symbolic logic was considered to be the core of artificial intelligence and linguistic theory (jones1994natural) and the most established textbooks on mathematical and computational linguistics from this era focus on discrete symbolisms, barely mentioning statistics, probability, or continuous mathematical foundations such as vectors and gradients (partee1993mathematical; allen1995natural).
@@
Today’s chatbots come from the very techniques that were supposed to have been left behind: machine learning in general and neural networks in particular are the cornerstones of AI today. During the 1980s, researchers including Hinton and LeCun developed the backpropagation method, which enables deep neural networks to be trained efficiently, by passing feedback on classification results back through layers in the network, using the chain rule from multivariable calculus (rumelhart1986learning). Neural nets built in this way started to outperform expert systems at several tasks, for example image recognition for handwritten digits (lecun1995convolutional). In 1997, Deep Blue, a chess computer program invented by Hsu, a doctoral student at Carnegie Mellon University, defeated the world chess champion Gary Kasparov, relying on a massively parallel expert fuzzy-rule-based system and effective use of a dataset of Grandmaster games (campbell2002deep). In 2000s, the first neural net model for next-word prediction was trained successfully to produce relatively consistent natural language phrases, over longer context distances than those achieved with $n$-gram models. AI was beginning to produce systems that outperformed humans on complex symbolic tasks and, rather than doing this using rules created by human domain experts, this was achieved by leveraging fast massive computation. The game of Go, more combinatorically demanding than chess by a wide margin, was mastered by the Alpha Go system, using databases of real games for training, and then even more strikingly, by Alpha Go Zero, which learned solely by playing against itself, with no human supervision except for being given the rules and the resources to simulate play (silver2017mastering).
@@
Early on, attempts at AI development focused on creating systems which would excel at specific tasks, assuming that such systems put together would create a human-like, general intelligence: for example, Shannon famously proposed that the game of chess would be an ideal challenge for developing intelligent systems, since it is generally considered to require “thinking” for skilful play (shannon1950chess). Based on the combinatoric explosion of possible chess positions, Shannon argued that a brute-force approach, leveraging large computational power, would at best enable the creation of a weak chess player. But, fast forward to a few decades later and it turns out that one can design an extremely effective chess system, mainly by applying brute-force computation over large datasets. This is both an insight and a challenge to humans; some of the most poignant (albeit wry) reflections on this theme have been written by Kasparov himself, the first human world chess champion to lose to a computer (or as (kasparov2017deep) reminds us, the last human to win a match against the world computer chess champion, a title he is likely to retain while records last).
@@
AI agents have become better by carefully leveraging larger example datasets and computational resources, a process which contrasts with the way humans accomplish similar expertise, typically requiring years of devotion (gladwell2008outliers). Understandably, this is often unwelcome news to human experts, and the theme that general methods that leverage computation are ultimately the most effective, and by a large margin has even been dubbed “The Bitter Lesson” (sutton2019bitter). But, the two questions, are we trying to construct systems which behave intelligently and are we trying to re-create human intelligence, are not the same, though the recognition of this difference has taken some time (russell2022artificial, Ch 1).
@@
The large language models (LLMs) that drive modern chatbots are a striking example of this engineering approach. They are built using neural networks to predict sequences of word-embedding vectors, the most recent key ingredient being the use of attention mechanisms, which enable sequential behaviour to be modelled using parallel computation (vaswani2017attention). Subsequent models, including GPT-3 (brown2020language), Llama (touvron2023llama), and DeepSeek (liu2024deepseek), are trained and optimized to whole new levels of performance, though the core components of word embeddings, feed-forward neural networks, and attention-heads using projection (with the whole system trained to optimize a loss-function throughout the network using backpropagation) are fundamentally those established in the transformer model of vaswani2017attention.
@@
These models use none of the syntactic parsers, logical forms, or knowledge representation techniques that dominated much of late-1900s AI. Instead, they work much as (turing1950computing) suggested: a ‘child machine’ is initialized with enormous computational resources and relatively few built-in procedures, just enough to learn from masses of training examples, after which it performs well — so well that its output, as exemplified by modern chatbots, produces language successfully enough that it would be considered a clear sign of intelligence in many established tests.